Compound Semi Annually.
P dollars is invested at annual interest rate r for 1 year. If the interest rate is compounded semiannually, then the polynomial P(1+r/2)squared presnts the value of the investment after 1 year. Rewrite this expression without parenthesis. Evalate the polynomial if P=$200 and r=10%
I need help understanding this type of word problem.

Answer by stanbon(57967)

P dollars is invested at annual interest rate r for 1 year. If the interest rate is compounded semiannually, then the polynomial P(1+r/2)squared presnts the value of the investment after 1 year. Rewrite this expression without parenthesis. Evalate the polynomial if P=$200 and r=10%
---------
A(t) = P(1+(r/n))^(nt)
Your Problem:
Amount invested is P
n = 2 meaning you compound the money twice in a year
t is the number of years; in your case that is "1"
r is the annual rate of interest
--------------
A(1) = P(1+(0.10/2))^(2*1)
A(1) = P(1.05)^2
A(1) = P(1.1025)
-----------
If you expand the expression as you were directed you get:
A(t) = P(1+(r/n))^2
= P[(1+(r/n))(1+(r/n))]
= P[1 + 2(r/n) + (r/n)^2]
----------------
If P=$200 and r=10% you get:
A(1) = 200[1 + 2*0.05 + (0.05)^2]
A(1) = $220.50
------------
Cheers,
Stan H.